# Ab initio spin Hamiltonians and magnetism of Ce and Yb triangular-lattice compounds

### v1

@@ Abstract
Low-energy Hamiltonians for those rare-earth triangular materials remain to date quite uncertain, in particular, due to lack of single crystals for delafossites [33]. It appears to be difficult to precisely extract small two-site anisotropies from magnon dispersions in an ordered state induced by applied field, hence, isotropic NN and NNN interactions are often assumed in the analysis [36] possible admixture of excited CF levels by magnetic field further complicating the analysis. Theoretically, ab initio calculations of exchange interactions in Yb delafossites on the basis of DFT electronic structure have so far predicted too small NNN couplings compared to experimental expectations [37, 36].
KCeO2 $0.85|5/2;\mp 1/2\rangle\pm 0.47|5/2;\pm 5/2\rangle\pm 0.18|7/2;\mp 1/2\rangle$
KCeS2 $0.90|5/2;\mp 1/2\rangle\pm 0.44|5/2;\pm 5/2\rangle-0.05|7/2;\mp 7/2\rangle$
RbCeO2 $0.87|5/2;\mp 1/2\rangle\pm 0.44|5/2;\pm 5/2\rangle\pm 0.17|7/2;\mp 1/2\rangle$
YZGO $0.73|7/2;\mp 7/2\rangle-0.50|7/2;\pm 5/2\rangle\mp 0.47|7/2;\mp 1/2\rangle$
Table 1: Ground-state Kramers doublet wavefunctions in the total-moment eigenstates $|JM\rangle$ basis. In the case of Ce oxides, the admixture of excited $J=7/2$ SO multiplet to the GS WV involve additional contributions apart from the largest one shown, as expressed by the dots at the end (see the SM [56] for the complete expressions).
@@
In this work, we tackle the problem of evaluating realistic spin Hamiltonians for triangular rare-earth systems using DFT+dynamical mean-field theory [38, 39, 40] treating strongly localized rare-earth 4ff states in a quasiatomic (Hubbard-I) approximation [41] to obtain the high-temperature paramagnetic electronic structure, as well as, the 4ff-shell CF splitting and ground-state Kramers doublet. Intersite exchange interactions (IEI) are evaluated from the same paramagnetic electronic structure using the force-theorem in Hubbard-I (FT-HI) method [42]. In contrast to the previous ab initio calculations deriving IEI in triangular rare-earth systems from DFT electronic structure with metallic 4ff bands [37], within our framework the IEI are obtained from realistic self-consistent DFT+HI electronic structure of the target compounds with localized 4ff states. A significant impact of DFT+DMFT charge-density self-consistency on the electronic structure and CF splitting is well documented [43, 44, 45]. We apply this framework to derive realistic spin Hamiltonians for a representative set of Ce- and Yb- triangular compounds, namely: KCeO2, KCeS2, RbCeO2, and YbZn2GaO5 (YZGO). These materials are either experimentally inferred to exhibit conventional magnetic orders (KCeO2 [23] and KCeS2 [22, 24]), a putative DSL (YZGO [33, 46]) or remain to be investigated in detail (RbCeO2, for which no order was found in a preliminary study [25], but a full characterization has not been carried out). We then study the resulting Hamiltonians through exact diagonalization (ED) calculations [47, 48], obtaining numerically exact ground- and low-energy eigenstates resolved by the lattice space group on clusters with sizes up to $36$ spins. This, then, allows us to analyze in detail the nature of the magnetic ground-state (GS) spin-spin correlations for each of the considered compounds.
@@
Ab initio method. We employ a charge self-consistent implementation [49, 50, 51] of DFT+DMFT based on the Wien2k linearized augmented-plane-wave (LAPW) full-potential code [52] and ”TRIQS” library [53]. All calculations were carried out using the reported experimental lattice structures [23, 24, 25, 33]. The IEI are calculated by FT-HI approach [42] using the MagInt code [54] analogously to previous applications of this approach to correlated magnetic insulators, see Ref. [55] for a review. The FT-HI method includes all kinetic exchange contributions due to virtual hopping of 4ff electrons. See Supplemental Material (SM) for more details of our ab initio calculations [56].
@@
Figure 1: Calculated CF level splitting. The thick red lines are calculated CF levels; the dashed blue lines are CF excitations measured in INS from Refs. [23, 22, 25, 33]. Note that three excitations are seen in INS spectra for the Ce compounds.
@@
Table 2: Calculated values for the in-plane ($g_{ab}$) and out-of-plane ($g_{c}$) gyromagnetic tensor components. Other values are obtained by electron spin resonance, Ref. [23] (a); quantum chemistry calculations, Ref. [22] (b), Ref. [24] (c), Ref. [25] (d); magnetization measurements, Ref. [33] (e).
@@
Crystal-field splitting and ground state. DFT+HI calculations converge to the expected insulating state with $f^{1}$ and $f^{13}$ GS occupancy of the Ce and Yb 4ff shells, respectively (see the SM for the calculated DFT+HI spectral functions). The average calculated SO coupling $\lambda$=87 meV for the Ce compounds leading to the SO gap between the GS $J=5/2$ and excited $J=7/2$ multiplets of 0.30 eV, with at most 1.5% spread among the 3 compounds. The value $\lambda=$0.39 eV with the SO splitting of 1.37 eV is obtained for Yb in YZGO. The calculated CF splitting are shown in Fig. 1 and the corresponding GS Kramers doublets are listed in Table 1 (see the SM for a complete list CF wavefunctions). We find a particularly large total CF splitting – about 60% of the SO gap – for the Ce oxides. Correspondingly, there is a non-negligible admixture $\approx 0.17|7/2;\pm 1/2\rangle$ of the SO excited multiplet to their GS Kramers doublets. The CF splitting in KCeS2 is almost 3 times smaller leading to the corresponding reduction of the $J=7/2$ admixture. No noticeable multiplet mixing is found for YZGO. One may notice the dominating contribution of $|M=\pm 1/2\rangle$ to the GS of the Ce systems corresponding to an in-plane magnetic anisotropy. In contrast, one may expect an out-of-plane anisotropy in YZGO due to the largest $|M=\pm 7/2\rangle$ contribution to its doublet. These expectations are confirmed by our results on the $g$-tensor (Table 2) extracted from the calculated magnetic moment operators $M_{\alpha}=g_{\alpha}S^{\alpha}$, $\alpha=x,y,z$ within the GS Kramers doublet. Here we introduced the pseudo-spin $S=1/2$ to label the states within the GS doublet and the corresponding spin operators $S^{\alpha}$. One sees that the in-plane components $g_{ab}=g_{x}(=g_{y})$ largely dominate in all Ce systems. In KCeS2, the out-of-plane component even becomes negative as a result of a near-perfect cancellation between the J=J=5/2 states, with the moment induced by a J=J=7/2 admixture. The magnitude of $g$ in YZGO is significantly larger, with a rather small out-of-plane anisotropy.
@@
The experimental determination of the CF splitting and ground state in the Ce compounds under consideration is complicated by the appearance of an additional mode in their inelastic neutron scattering spectra [21, 25, 22]. A $J=5/2$ multiplet split into three Kramers doublet should result in two measurable CF excitations; instead, three excitations are observed. The lowest SO excited $J=7/2$ level is predicted by our calculations to be at about 300 meV agreeing with experiment [23]. Secondary oxide phases [24], vibronic splitting [57, 58] or more complicated multisite many-body effects [59] have been considered as possible origins of the additional mode, which is also observed in layered Ce intermetallics [59] and Ce pyrochlores [60], but it remains so far an unsolved puzzle. Taking this complication into account, our predictions for the CF levels energies in all compounds under consideration are in good agreement with experiment (Fig. 1), albeit with some systematic underestimation of the splitting by the present theory. Similarly, a good agreement is found for the $g$-tensor (Table 2) in comparison with previous experimental estimates and quantum chemistry calculations. We note that no signal was observed in the electronic spin resonance for the out-of-plane direction in KCeS2 suggesting $g_{c}\ll g_{ab}$ [22].
@@
Effective magnetic Hamiltonian. Having obtained the GS Kramers doublet in all systems, we subsequently calculated the IEI within this manifold using the FT-HI method. Besides IEI, we also included the classical intersite dipole-dipole interaction, which can be important given rather small magnitudes of anisotropic IEI in the target materials, evaluated using the ab initio gyromagnetic tensors (Table 2).
@@
where $\mathbf{S}^{T}_{i}=[S^{x}_{i},S^{y}_{i},S^{z}_{i}]$, $\Delta$ is the diagonal XXZ anisotropy, $J_{\pm\pm}$ is the diagonal XY anisotropy and $J_{z\pm}$ are off-diagonal terms. The coupling for the NNN bond along $\mathbf{r}_{0}^{\mathrm{NNN}}=[010]||y$ takes the same form of Eq. (1); we will use the prime to distinguish the NNN terms ($^{\prime},J±±J_{\pm\pm}^{\prime}$ etc.). The coupling for other NN and NNN bonds follows from the symmetry of the relevant point group that is generated by the threefold rotation about the $z$ axis, $C_{3}$, two-fold rotation $C_{2}$ around each axis $\mathbf{a}_{1},\mathbf{a}_{2},\mathbf{a}_{3}$, of the triangular lattice, and inversion $I$. The spin operator vector $\mathbf{S}$ transforms as a usual 3D vector upon proper rotations [63, 62]. By applying the $C_{3}$ and $C_{3}^{2}$ rotations about the $z$ axis, one obtains the couplings for the corresponding bonds as $\hat{J}_{\mathbf{r}}=\mathcal{R}\hat{J}_{\mathbf{r}_{0}}\mathcal{R}^{T}$, where $\mathbf{r}=\mathcal{R}\mathbf{r}_{0}$ and $\mathcal{R}$ is the corresponding rotation matrix. The couplings for the bonds related by inversion are identical, since inversion does not affect the spins. The symmetry of Eq. (1) is not imposed in FT-HI calculations, in which all matrix elements of (1) are evaluated independently, nevertheless it appears in the calculated IEI.
@@
Figure 2: Spin structure factors and low-energy spectrum for the YZGO (a,b) and KCeS2 (c,d) compounds in a $36$-site cluster. The isotropic structure factor $\mathcal{S}\left(\bm{k}\right)$ in (a) is peaked at the $K$ points, indicating $120^{\circ}$ order, consistent with low-energy excitations at $K_{0}.\text{A}$ and $K_{1}.\text{B}$. The peak at $M_{2}$ of $\mathcal{S}^{yy}(\bm{k})$ in (c) indicates stripe-$\perp$ order, consistent with low-energy excitations at the $M$ points in (d).
To assess the overall agreement in magnitude with experiments, we calculated high-temperature susceptibility by solving (1) in mean-field in an applied field using the McPhase package [64] and an in-house module. To compare with polycrystalline data we averaged the calculated susceptibility as $(2\chi_{ab}+\chi_{c})/3$. For all oxide systems we find quite good agreement of the calculated Curie-Weiss temperatures $\theta$ with available experimental data (Table 3). In the case of KCeS2, theoretical intersite couplings are severely underestimated. This is also apparent from comparison of our result with the estimates of Ref. [65] that were obtained by quantum chemistry calculations and subsequently adjusted to agree with experimental INS. 4ff kinetic exchange captured by FT-HI does not, apparently, give the leading contribution to the IEI in this sulfide.
Low-energy spectra and ground-state correlations. We now investigate the GS phase diagram of the effective magnetic Hamiltonians introduced in Eq. (1) with ED using the XDiag library [48]. We obtain symmetry-resolved excitation spectra employing translational symmetries and the $I$ lattice inversion symmetry without spin rotation on cluster sizes $N=12,32,36$, see the SM [56] for the employed lattice geometries. The irreducible representations (irreps) are represented by $\bm{k}.\rho$, where $\bm{k}$ labels the momentum and $\rho$ labels the inversion irrep, i.e. either $\rho=\text{A}$ for the even irrep or $\rho=\text{B}$ for odd irrep. We note that the additional point group symmetries are not employed, as an implementation of combined lattice-spin rotations is not readily accessible. We also obtained the GS static spin structure factors,
@@
where $\alpha=x,y,z$, $N$ denotes the number of lattice sites and $\bm{r}_{n}$ the position of the $n^{\rm th}$ spin. Furthermore, we denote by,
$\mathcal{S}\left(\bm{k}\right)=\mathcal{S}^{xx}\left(\bm{k}\right)+\mathcal{S}^{yy}\left(\bm{k}\right)+\mathcal{S}^{zz}\left(\bm{k}\right),$ (3)
the isotropic spin structure factor.
@@
In panels $a)$ and $b)$ of Fig. 2, we show $\mathcal{S}\left(\bm{k}\right)$ and the low-lying energy spectrum for the model parameters of YZGO (see table 3) on the $N=36$ cluster. Both the organization of the spectrum and the spin structure factor are consistent with the system realizing the $120^{\circ}$ AFM phase. The ground state transforms under the $\Gamma.\text{A}$ irrep, while the first excited states—nearly degenerate—belong to the $\Gamma.\text{B}$ and $K.\text{A}$ irreps. These symmetry quantum numbers match the expected signatures of the $120^{\circ}$ AFM phase [66, 67], see the SM [56] for further details on how these quantum number are predicted. A key signature of a DSL is a gapless singlet mode at the $X_{0}$ point in the first Brillouin zone (FBZ), which constitutes the singlet monopole excitations [19]. However, the excitation energy at $X_{0}$ is significantly larger compared to the putative DSL of the $J_{1}$-$J_{2}$ model [11], rendering this scenario unlikely. Furthermore, the static spin structure factor exhibits a pronounced peak at crystal momentum $K$ within the FBZ, in agreement with the known ordering wave vector of the $120^{\circ}$ AFM state [68, 69]. Analogously, we observe spin correlations and a low-energy spectrum consistent with a $120^{\circ}$ state for the effective Hamiltonians of the KCeO2 and RbCeO2 compounds, see the SM [56].
@@
Figure 3: $a)$ and $d)$ - Approximate phase diagram of the effective magnetic Hamiltonian, inferred from the organization of quantum numbers in the low-energy spectrum. The dark red region indicates the $120^{\circ}$ antiferromagnetic phase, dark blue the stripy-$\parallel$ order, pale purple the stripy-$\perp$ order, and orange the regime where the first non-zero momentum excitation is in the $X_{1}.\text{A}$ irrep, which suggests a possible DSL phase. A typical classical ordering for each stripe phase is shown in the sketch inside panel d). The remaining panels show the static spin structure factor at selected momenta in the FBZ (see inset of panel $a$). Panels $b$ and $e$ display $\mathcal{S}(\mathbf{k})$ at $Z$, while $c$ and $f$ show the average $\tilde{\mathcal{S}}(M)$ over the three inequivalent $M$ points. In panel $b)$ we locate the parameters compatible with the YZGO system. Parameters: $\Delta=1.03J$, $J^{\prime}=0.02J$ in a)a)$c)$, and $J_{z\pm}=0$ in d)d)$f)$.
@@
The spectrum and correlations for the KCeS2 parameters, on the other hand, indicate a collinear stripy antiferromagnetic GS, as shown in panels $c)$ and $d)$ of Fig. 2. The low energy spectra reveals an approximate GS degeneracy between the $\Gamma$.A, and $M.\text{A}$ irreps, consistent with the stripy Néel state [67]. Accordingly, we observe a pronounced peak in $\mathcal{S}^{yy}(\bm{k})$ and $\mathcal{S}^{zz}(\bm{k})$ at momentum $M_{2}$We observe the peak only at one of the nonequivalent $M$ wave vectors, since the $C_{3}$ rotational symmetry of the Hamiltonian is slightly broken by numerical rounding in the coupling constants obtained from the ab-initio calculations.. As noticed previously [71], the symmetries of Eq. 1 allow for two distinct stripe orders, differing by how they transform under the combined lattice-spin rotation around the triangular lattice axis directions. The first, we here call stripe-$\parallel$ (stripe-parallel), aligns the spins along the direction of the stripe. It is invariant under the $C_{2}$ axis rotations as the spins align with the axis of rotation. The stripe-$\perp$ (stripe-perpendicular) state aligns the spins perpendicular to the stripe direction, and thus transforms non-trivially under the $C_{2}$ axis rotation. A representative configuration of each stripe type is shown in Fig. 3d). In general, this state is not invariant under the $C_{2}$ axis rotation. Thus, the (in)variance under this symmetry constitutes the distinguishing feature between the two states. We prefer to refer to these states as stripe-$\parallel$ and stripe-$\perp$ as we think this more precisely captures the essence of these phases than their previously used names stripe-$x$ and stripe-$yz$, respectively.
@@
As the ground state stripe stabilized is oriented along the $x$-direction of the lattice, the peak in the $\mathcal{S}^{yy}(\bm{k})$ and $\mathcal{S}^{zz}(\bm{k})$ at $M_{2}$ indicates the realization of the stripe-$\perp$ state, see the SM for further details. Remarkably, this is in full agreement with experiments on KCeS2 [24].
@@
We obtain an approximate phase diagram, shown in panel $a)$ and $d)$ of Fig. 3, based on the quantum numbers of the spectrum, which we show in the supplementary material. We begin by fixing $J^{\prime}=0.02J$ (a value numerically compatible with the largest matrix elements of the NNN coupling $\hat{J^{\prime}}$ in YZGO; see Table 3) and study the phase diagram as a function of the diagonal XY anisotropy $J_{\pm\pm}$ and the off-diagonal XY coupling $J_{z\pm}$. This corresponds to the panels (a,b,c) in the upper row of Fig. 3. In this case, the phase diagram consists of three distinct magnetically ordered states, namely, a $120^{\circ}$ AFM region, a stripe-$\parallel$ phase, and a stripe-$\perp$ phase, which correspond to the dark red, dark blue, and pale purple regions shown in panel $a)$, respectively.
@@
As on the 32-site cluster, the $K$ momentum is not resolved, the structure factor peaks at nearby wave vectors $Z_{0}$ and $Z_{1}$ in the $120^{\circ}$ AFM phase, as shown in panel (b) of Fig. 3. This reflects the frustrated realization of $120^{\circ}$ AFM order in this cluster. In contrast, the 12-site cluster resolves the $K$ points, showing a structure factor peak there consistent with the $120^{\circ}$ AFM order (see supplementary material). We observe in panel b) that the couplings determined for the YZGO system are within the $120^{\circ}$ AFM as expected.

### v2

@@ Abstract
Low-energy Hamiltonians for those rare-earth triangular materials remain to date quite uncertain. In particular, due to lack of single crystals for majority of delafossites [33], two-site anisotropies were assessed experimentally only in few cases [36, 27]. Hence, isotropic NN and NNN interactions are often assumed in the analysis [37], which is further complicated by possible admixture of excited CF levels by magnetic field. Theoretically, ab initio calculations of exchange interactions in Yb delafossites on the basis of DFT electronic structure have so far predicted too small NNN couplings compared to experimental expectations [38, 37].
@@
In this work, we tackle the problem of evaluating realistic spin Hamiltonians for triangular rare-earth systems using DFT+dynamical mean-field theory (DMFT) [39, 40, 41] treating strongly localized rare-earth 4ff states in a quasiatomic (Hubbard-I) approximation [42] to obtain the high-temperature paramagnetic electronic structure, as well as, the 4ff-shell CF splitting and ground-state Kramers doublet. Intersite exchange interactions (IEI) are evaluated from the same paramagnetic electronic structure using the force-theorem in Hubbard-I (FT-HI) method [43]. In contrast to the previous ab initio calculations deriving IEI in triangular rare-earth systems from DFT electronic structure with metallic 4ff bands [38], within our framework the IEI are obtained from realistic self-consistent DFT+HI electronic structure of the target compounds with localized 4ff states. A significant impact of DFT+DMFT charge-density self-consistency on the electronic structure and CF splitting is well documented [44, 45, 46]. We apply this framework to derive realistic spin Hamiltonians for a representative set of Ce- and Yb- triangular compounds, namely: KCeO2, KCeS2, RbCeO2, and YbZn2GaO5 (YZGO). These materials are either experimentally inferred to exhibit conventional magnetic orders (KCeO2 [23] and KCeS2 [22, 24]), a putative DSL (YZGO [33, 47]) or remain to be investigated in detail (RbCeO2, for which no order was found in a preliminary study [25], but a full characterization has not been carried out). We then study the resulting Hamiltonians through exact diagonalization (ED) calculations [48, 49], obtaining numerically exact ground- and low-energy eigenstates resolved by the lattice space group on clusters with sizes up to $36$ spins. This, then, allows us to analyze in detail the nature of the magnetic ground-state (GS) spin-spin correlations for each of the considered compounds. Previous works have already investigated certain regions of the parameter space for these type of Hamiltonians, through classical Monte Carlo studies [50, 51], ED [52] and notably using DMRG [32, 53, 54, 55].
@@
Ab initio method. We employ a charge self-consistent implementation [56, 57, 58] of DFT+DMFT based on the Wien2k linearized augmented-plane-wave (LAPW) full-potential code [59] and ”TRIQS” library [60]. All calculations were carried out using the reported experimental lattice structures [23, 24, 25, 33]. The IEI are calculated by FT-HI approach [43] using the MagInt code [61] analogously to previous applications of this approach to correlated magnetic insulators, see Ref. [62] for a review. The FT-HI method includes all kinetic exchange contributions due to virtual hopping of 4ff electrons. See the Supplemental Material (SM) [63] for more details on our ab initio calculations and the effect of charge-density self-consistency in DFT+HI.
@@
Figure 1: Calculated CF level splitting. The thick red lines are calculated CF levels; the dashed blue lines are CF excitations measured in inelastic neutron scattering (INS) from Refs. [23, 22, 25, 33]. Note that three excitations are seen in INS spectra for the Ce compounds.
@@
Table 1: Calculated values for the in-plane ($g_{ab}$) and out-of-plane ($g_{c}$) gyromagnetic tensor components. Other values are obtained by electron spin resonance (ESR), Refs. [23] (a) ESR and quantum chemistry calculations (QCC) [22] (b); QCC, Ref. [24] (c); INS, Ref. [25] (d); magnetization measurements, Ref. [33] (e).
@@
Crystal-field splitting and ground state. DFT+HI calculations converge to the expected insulating state with $f^{1}$ and $f^{13}$ GS occupancy of the Ce and Yb 4ff shells, respectively (see the SM for the calculated DFT+HI spectral functions). The calculated CF splitting are shown in Fig. 1 and the corresponding Kramers doublets are listed in the SM [63]. We find a particularly large total CF splitting – about 60% of the SO gap (the latter is 0.3 eV in all Ce systems) – for the Ce oxides. Correspondingly, there is a non-negligible admixture $\approx 0.17|7/2;\pm 1/2\rangle$ of the SO excited multiplet to their GS Kramers doublets. The CF splitting in KCeS2 is almost 3 times smaller leading to the corresponding reduction of the $J=7/2$ admixture. No noticeable multiplet mixing is found for YZGO due to its much larger SO gap of 1.37 eV. The dominating contribution of $|M=\pm 1/2\rangle$ to the GS of all Ce systems leads to an in-plane magnetic anisotropy. In contrast, the $|M=\pm 7/2\rangle$ states provide the largest contribution to the GS doublet of YZGO. The corresponding $g$-tensor extracted from the calculated magnetic moment operators $M_{\alpha}=g_{\alpha}S^{\alpha}$, $\alpha=x,y,z$ within the GS Kramers doublet is listed in Table 1. Here we introduced the pseudo-spin $S=1/2$ to label the states within the GS doublet and the corresponding spin operators $S^{\alpha}$. One sees that the in-plane components $g_{ab}=g_{x}(=g_{y})$ largely dominate in all Ce systems. In KCeS2, the out-of-plane component even becomes negative as a result of a near-perfect cancellation between the J=J=5/2 states, with the moment induced by a J=J=7/2 admixture. The $g$-tensor in YZGO is significantly larger and quasi-isotropic.
@@
The experimental determination of the CF splitting and GS in the Ce compounds under consideration is complicated by the appearance of an additional, third, mode in their INS spectra [21, 25, 22], though only two CF excitations are expected in the case of a $J=5/2$ GS. This phenomenon, observed also in other Ce systems, remains so far without commonly accepted explanation, see Refs. [24, 64, 65]. Taking this complication into account, our predictions for the CF levels energies in all compounds under consideration are in good agreement with experiment (Fig. 1), albeit with some systematic underestimation of the splitting by the present theory. Similarly, a good agreement is found for the $g$-tensor (Table 1) in comparison with previous experimental estimates and quantum chemistry calculations. We note that no signal was observed in the electronic spin resonance for the out-of-plane direction in KCeS2 suggesting $g_{c}\ll g_{ab}$ [22].
@@
Effective magnetic Hamiltonian. The IEI within the GS Kramers manifold were calculated by the FT-HI method. Besides IEI, we also included the classical intersite dipole-dipole interaction, which can be important given rather small magnitudes of anisotropic IEI in the target materials, evaluated using the ab initio gyromagnetic tensors (Table 1).
@@
where $\mathbf{S}^{T}_{i}=[S^{x}_{i},S^{y}_{i},S^{z}_{i}]$, $\Delta$ is the diagonal XXZ anisotropy, $J_{\pm\pm}$ is the diagonal XY anisotropy and $J_{z\pm}$ are off-diagonal terms. The coupling for the NNN bond along $\mathbf{r}_{0}^{\mathrm{NNN}}=[010]||y$ takes the same form of Eq. (1); we will use the prime to distinguish the NNN terms ($^{\prime},J±±J_{\pm\pm}^{\prime}$ etc.). The coupling for other NN and NNN bonds follows from the symmetry [67, 66], see SM [63] for details.
@@
Figure 2: Spin structure factors and low-energy spectrum for the YZGO ($a,b$) and KCeS2 ($c,d$) compounds in a $36$-site cluster. The isotropic structure factor $\mathcal{S}\left(\bm{k}\right)$ in $a)$ is peaked at the $K$ points, indicating $120^{\circ}$ order, consistent with low-energy excitations at $K_{0}.\text{A}$ and $K_{1}.\text{B}$. The peak at $M_{2}$ of $\mathcal{S}^{yy}(\bm{k})$ in $c)$ indicates stripe-$\perp$ order, consistent with low-energy excitations at the $M$ points in $d)$.
To assess the overall agreement in magnitude with experiments, we calculated high-temperature susceptibility by solving (1) in mean-field in an applied field using the McPhase package [68] and an in-house module. For all oxide systems we find quite good agreement of the calculated Curie-Weiss temperatures $\theta$, extracted from the averaged susceptibility $(2\chi_{ab}+\chi_{c})/3$, with available experimental data (Table 2). In the case of KCeS2, theoretical intersite couplings are severely underestimated. This is also apparent from comparison of our result with the estimates of Ref. [69] that were obtained by quantum chemistry calculations and subsequently adjusted to agree with experimental INS. 4ff kinetic exchange captured by FT-HI does not, apparently, give the leading contribution to the IEI in this sulfide.
Low-energy spectra and ground-state correlations. We now investigate the GS phase diagram of the effective magnetic Hamiltonians introduced in Eq. (1) with ED using the XDiag library [49, 70]. We obtain symmetry-resolved excitation spectra employing translational symmetries and the lattice inversion symmetry without spin rotation on cluster sizes $N=12,32,36$, see the SM [63] for the employed lattice geometries. The irreducible representations (irreps) are represented by $\bm{k}.\rho$, where $\bm{k}$ labels the momentum and $\rho$ labels the inversion irrep, i.e. either $\rho=\text{A}$ for the even irrep or $\rho=\text{B}$ for odd irrep. We note that the additional point group symmetries are not employed, as an implementation of combined lattice-spin rotations is not readily accessible. We also obtained the GS static spin structure factors,
@@
where $\alpha=x,y,z$, $N$ denotes the number of lattice sites and $\bm{r}_{n}$ the position of the $n^{\rm th}$ spin. Furthermore, we denote by $\mathcal{S}\left(\bm{k}\right)=\sum_{\alpha}\mathcal{S}^{\alpha\alpha}\left(\bm{k}\right)$, the isotropic spin structure factor.
@@
In panels $a)$ and $b)$ of Fig. 2, we show $\mathcal{S}\left(\bm{k}\right)$ and the low-lying energy spectrum for the model parameters of YZGO (see table 2) on the $N=36$ cluster. Both the organization of the spectrum and the spin structure factor are consistent with the system realizing the $120^{\circ}$ AFM phase. The GS transforms under the $\Gamma.\text{A}$ irrep, while the first excited states (nearly degenerate) belong to the $\Gamma.\text{B}$ and $K.\text{A}$ irreps. These symmetry quantum numbers match the expected signatures of the $120^{\circ}$ AFM phase [71, 72], see the SM [63] for further details on how these quantum numbers are predicted.
A key signature of a DSL is a gapless singlet mode at the $X_{0}$ point in the first Brillouin zone (FBZ), corresponding to the singlet monopole excitations [19]. Although in the present case we cannot identify these excitations as singlets due to the absence of spin SU(2) symmetry, they still occur at the same momentum. In contrast, the excitation energy at $X_{0}$ is significantly larger compared to the putative DSL of the $J_{1}$-$J_{2}$ model [11], rendering this scenario unlikely. Furthermore, $\mathcal{S}(\bm{k})$ exhibits a pronounced peak at crystal momentum $K$, in agreement with the known ordering wave vector of the $120^{\circ}$ AFM state [2, 73]. Analogously, we also observe spin correlations and a low-energy spectrum consistent with a $120^{\circ}$ state for the effective Hamiltonians of the KCeO2 and RbCeO2 compounds, see the SM [63].
@@
Figure 3: $a)$ and $d)$ - Approximate phase diagram of the effective magnetic Hamiltonian, inferred from the quantum number organization of the low-energy spectrum. Dark red marks the $120^{\circ}$ AFM phase, dark blue the stripy-$\parallel$ order, pale purple the stripy-$\perp$ order, and orange the regime with first non-zero momentum excitation in the $X_{1}.\text{A}$ irrep, suggesting a possible DSL phase. A typical classical ordering for each stripe phase is shown in the sketch in $d)$. The remaining panels show the static spin structure factor at selected momenta in the FBZ (see inset in $a$)): panel $b)$ and $e)$ at $Z$, while $c)$ and $f)$ show the average over the three inequivalent $M$ points. In $b)$ we identify the parameters compatible with the YZGO system. Parameters: $\Delta=1.03J$, $J^{\prime}=0.02J$ in a)a)$c)$, and $J_{z\pm}=0$ in d)d)$f)$.
@@
On the other hand, the spectrum and correlations for the KCeS2 system indicate a collinear stripy antiferromagnetic GS, as shown in panels $c)$ and $d)$ of Fig. 2. The low energy spectra reveals an approximate GS degeneracy between the $\Gamma$.A, and $M.\text{A}$ irreps, consistent with the stripy AFM state [72]. Accordingly, we observe a pronounced peak in $\mathcal{S}^{yy}(\bm{k})$ and $\mathcal{S}^{zz}(\bm{k})$ at momentum $M_{2}$The peak appears at only one $M$ point because slight numerical rounding in the ab-initio couplings weakly breaks the Hamiltonian’s $C_{3}$ symmetry.. As noted [50], the symmetries of Eq. (1) allow for two distinct stripe orders, differing by how they transform under the combined lattice-spin rotation around the triangular lattice axis directions. The first aligns the spins along the direction of the stripe, we call it stripe-$\parallel$ (stripe-parallel) (this is also named stripe-$x$). It is invariant under the $C_{2}$ axis rotations, as the spins align with it. In the second, the spins align perpendicular to the stripe direction, and thus transforms non-trivially under the $C_{2}$ axis rotation. We name it stripe-$\perp$ (stripe-perpendicular) (also known as stripe-$yz$). Thus, the (in)variance under this symmetry constitutes the distinguishing feature between the two states . A representative configuration of each stripe type is shown in Fig. 3 $d)$.
@@
The stabilized stripe is oriented along the $x$-direction of the lattice, so, the peak in the $\mathcal{S}^{yy}(\bm{k})$ at $M_{2}$ indicates the realization of the stripe-$\perp$ state, see the SM for further details. Remarkably, this is in complete agreement with experiments on KCeS2 [24].
@@
We perform extensive ED calculations on the $32$-site cluster and use the $12$-site cluster to corroborate the phase diagram obtained, whose results and discussion are left to the SM [63].
@@
We obtain an approximate phase diagram (Fig. 3 $a)$ and $d)$) based on the quantum numbers of the spectrum (check the SM). We begin by fixing $J^{\prime}=0.02J$ (a value compatible with the largest matrix elements of the NNN coupling $\hat{J^{\prime}}$ in YZGO; see Table 2) and study the phase diagram as a function $J_{\pm\pm}$ and $J_{z\pm}$. This corresponds to the panels ($a,b,c$) in the upper row of Fig. 3. In this case, the phase diagram consists of three distinct magnetically ordered states, a $120^{\circ}$ AFM region, a stripe-$\parallel$ phase, and a stripe-$\perp$ phase, which correspond to the dark red, dark blue, and pale purple regions shown in panel $a)$, respectively.
@@
As in the 32-site cluster, the $K$ momentum is not resolved, $\mathcal{S}(\bm{k})$ peaks at nearby wave vectors $Z_{0}$ and $Z_{1}$ in the $120^{\circ}$ AFM phase, as shown in Fig. 3 $b)$. This reflects the frustrated realization of $120^{\circ}$ AFM order in this cluster.